OpenStudy (anonymous):

Find the length of the side of a square whose diagonals are 20.

OpenStudy (anonymous):

 now ABC is right triangle and as it is square so x=y use Pythagoras theorem for right triangle you'll get.. x^2 + y^2 = 20^2 as x=y so you can write x^2 + x^2 = 400 (replaced "y" with "x" as both are same) 2x^2 = 400 x^2 = 400/2 x^2 = 200 x = sqrt (200)